10 questions? PLEASE HELP ME SOLVE THESE. THANKYOU.
Instructions: Solve each equation. Change fractions to decimals. Enter only the numerical answer in the blank.
1. 7x+3=15x+9
Answer:
2. (1/2)=(2/3)b+(1/6)b
Answer:
3. 0.3a+0.7=0.5a-0.1
Answer:
4. 2n-5=8n+7
Answer:
5. 3+4x=3x+6
Answer:
6. 5(2w-4)=6w
Answer:
7. FOR THE FOLLOWING PROBLEMS, DO NOT INCLUDE UNITS YOUR ANSWERS.
Suppose you hike up a hill at 4km/h. You hike down at 6km/h. Your hiking trip took 3 hours. How much did it take you to hike up the hill. ENTER YOUR DECIMAL AS A DECIMAL.
8. The length of a rectangle is 8in. More than its width. The perimeter of the rectangle is 24in. What is the width of the rectangle?
9. The sum of two consequative EVEN integers is 118. what is the LESSER of the two integers.
10. You and your brother started biking at noon from places that are 52 mi apart. You rode toward eachother and met at 2:00PM. Your brothers average speed was 4mi/h faster than your average speed. What is your BROTHERS speed?

Thankyou guys for your help! I hope you get to this fast,

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

1. 7x+3=15x+9
Subtract 7x
3 = 8x + 9
Subtract 9
-6 = 8x
8x = -6
x = -6/8 = -3/4
-------------
2. (1/2)=(2/3)b+(1/6)b
Multiply thru by 12
6 = 8b + 2b
Similar to #1
-------------
3. 0.3a+0.7=0.5a-0.1
Multiply by 10
3a + 7 = 5a - 1
Similar to #1
----------------
4. 2n-5=8n+7
Do like #1
---------------
5. 3+4x=3x+6
More of the same.

Answer by ewatrrr(24785)


Hi, recommend showing what work you have done and posting only 1 or 2 two questions that are similar in content.
"I hope You get to this fast" ...must admit, did put a smile in my heart. :)
1. 7x+3=15x+9 | 8x = -6 |finish to decimal answer
2. (1/2)=(2/3)b+(1/6)b | b = (6/5)*(1/2) |finish to decimal answer
3. 0.3a+0.7=0.5a-0.1 | .8/.2 = a |finish to decimal answer
4. 2n-5=8n+7 | 6n = -12 |finish it up
5. 3+4x=3x+6 |x = 3
6. 5(2w-4)=6w | 4w = 20 |finish it up
7. hike up a hill at 4km/hr.. hike down at 6km/hr. hiking trip took 3 hours.
4kmp*t = 6kmph*(3hr-t) | 10t = 18 |finish it up, t is time uphill
8. length of a rectangle is 8in. More than its width. P = 24in
2x + 2(x+8) = 24in | 4x = 8in |finish it up, x is the width
9. The sum of two consecutive EVEN integers is 118
x + (x+2) = 118 | 2x = 116 |finish it up, x is the lesser integer
10. Your brothers average speed was 4mi/h faster than your average speed.
together they rode 52mi towards each other and met in 2hrs.
2x + 2(x-4mph) = 52mi | 4x = 60 |finish it up, x is the brother's speed